# Fine-Grained Image-Text Alignment in Medical Imaging Enables Cyclic Image-Report Generation

### v4

@@ 3.2 AdaMatch-based LLM for Cyclic CXR-Report Generation (AdaMatch-Cyclic)
Figure 3: Instruction data example for CXR-to-report generation.
@@ 3.2.1 CXR-to-Report Generation
In CXR-to-report generation, we employ the AdaMatch to match a predefined textual codebook with the CXR image to obtain the matched keywords as hints to guide LLM to generate reports.
Building textual codebook. Specifically, to establish a textual codebook, we use a pre-trained BioEN [35] to extract the related entities from the medical reports in the training set, where the related entities are divided into four entity groups, i.e., biological structure, detailed description, disease disorder, and sign symptom. Next, we compute the frequency of each entity and pick top $\kappa_{0}$ entities for each entity group as keywords in the textual codebook.
Keywords Extraction. With the textual codebook, we employ the frozen AdaMatch model to match the keywords from the textual codebook with the adaptive patches of CXR images, and obtain a patch-word similarity matrix $s^{P}\in\mathbb{R}^{M\times N}$ between the keyword tokens and adaptive patch embeddings, where $M$ and $N$ denote the number of keyword tokens and adaptive patches. To extract the most matched keywords for each adaptive patch, we rank the patch-word similarity along the dimension of keyword tokens, obtain the top $\kappa_{1}$ patch-word similarities for each adaptive patch, and extract the corresponding keywords $k^{T}$. The matched keywords can explain potential lesions in each adaptive patch, to assist LLM in generating medical reports.
Figure 4: Instruction data example for report-to-CXR generation.
Instruction tuning LLM. After keywords extraction, we use a frozen VQ-GAN [13] encoder $E$ to encode the input CXR image $x^{I}$ to the quantized image latent vectors as image tokens $E(x^{I})$. For CXR-to-report generation, the dolly-v2-3b [9] model is adopted as the bidirectional LLM, which is pre-trained on the instruction-following dataset. When fine-tuning LLM, we convert CXR-to-report generation dataset in instruction-following format. An example of instruction data for CXR-to-report generation is shown in Fig. 3, including the instruction, input, keywords, and response parts, where input is the image tokens of the input CXR image, keywords are extracted by AdaMatch, and response is the ground-truth medical report. During training, the response part is hidden and the LLM learns to generate it in autoregressive manner.
By adopting AdaMatch-Cyclic for CXR-to-report generation, we can locate the potential lesion in CXR images, to provide more explainability for the generation process.
@@ 3.2.2 Report-to-CXR Generation
In report-to-CXR generation, we leverage the AdaMatch to match a predefined visual codebook with the generated reports $x^{T^{\prime}}$ to acquire matched keypatches as guidance for LLM to synthesize CXR images, where ‘keypatches’ are important image patches related to generated reports.
Building visual codebook. Concretely, we first construct a visual codebook with the most common adaptive patches as keypatches. To collect the most common adaptive patches for CXR images in the training set, the AdaMatch is utilized to match the adaptive patches of CXR images with textual tokens of medical reports to obtain top $\kappa_{2}$ CXR-report pairs with the highest report-to-CXR similarities $s^{T}$. For each CXR-report pair, we compute the word-patch maximum similarity $\max_{(1\leq n\leq N)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$ for each textual token, rank the word-patch maximum similarities, and extract the adaptive patches for top $\kappa_{3}$ similarities as keypatches in the visual codebook. Each keypatch includes its adaptive patches and the corresponding features.
Keypatches Extraction. With the visual codebook, the frozen AdaMatch matches the features of keypatches in the visual codebook with textual tokens of the generated report to acquire the word-patch similarity matrix $s^{W}\in\mathbb{R}^{(\kappa_{2}\times\kappa_{3})\times K}$, where $K$ denotes the number of textual tokens for the generated report. To obtain keypatches related to the generated report, we rank the word-patch similarity along the dimension of keypatches, obtain the top $\kappa_{4}$ word-patch similarity for each textual token, and extract the features of corresponding keypatches $k^{I}$.
Instruction tuning LLM. After keypatches extraction, we employ a frozen VQ-GAN encoder to convert the matched keypatches $k^{I}$ into image tokens $E(k^{I})$, and feed the pre-trained LLM with the instruction, the generated report, and image tokens of keypatches in the instruction-following format, as shown in Fig. 4. Then, LLM predicts the image tokens and the VQ-GAN decoder decodes the image tokens into the corresponding generated CXR image $x^{I^{\prime}}$.
@@ 3.2.3 Overall Objective
To optimize AdaMatch-Cyclic, we apply the standard language modeling objective for both CXR-to-report and report-to-CXR generation. For CXR-to-report generation, LLM is fed with the instruction $I^{t}$, image tokens of input CXR image $E(x^{I})$, and matched keywords $k^{T}$ in instruction-following format, and LLM aims to generate the ground-truth medical report of response part autoregressively. We compute the $P(u_{k})$ conditional probability of the $k$-th token, $P(u_{k})=\mathcal{LLM}(I^{t},E(x^{I}),k^{T})$, where $k$ is the token index after the response key (### Response:). The report generation loss $L_{text}$ for the response area is calculated by,
$L_{text}=\sum_{i=k}^{n}-\log P(u_{i}|u_{1},u_{2},\cdots,u_{i-1}),$ (10)
where $[u_{1},u_{2},\cdots,u_{i-1}]$ denotes the tokenized texts before the response part, and $n$ represents the maximum length of output tokens.
Figure 5: Qualitative comparison with existing methods in CXR-to-report generation on the MIMIC-CXR (1st row) and OpenI datasets (2nd row). The texts in different colors show similar meanings. The keywords on the right are obtained from AdaMatch model.
Similarly, for report-to-CXR generation, we feed the LLM with the instruction $I^{v}$, the generated report $x^{T^{\prime}}$, and the image tokens of keypatches $E(k^{I})$ in the instruction-following format, and obtain the conditional probability of $k$-th token $w_{k}$, $P(w_{k})=\mathcal{LLM}(I^{v},x^{T^{\prime}},E(k^{I}))$. The CXR generation loss $L_{img}$ is defined as:
$L_{img}=\sum_{i=k}^{n}-\log P(w_{i}|w_{1},w_{2},\cdots,w_{i-1}),$ (11)
With $L_{text}$ and $L_{img}$, the LLM can implement the bidirectional CXR and report generation according to instructions.
Table 1: CXR-to-report generation performance in comparison to existing methods on MIMIC-CXR and OpenI datasets.
MIMIC-CXR OpenI
Methods BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L
R2Gen [7] (2020) 0.3553 0.2232 0.1523 0.1038 0.1412 0.2784 0.3992 0.2407 0.1518 0.0973 0.1390 0.3052
R2GenCMN [8] (2021) 0.3719 0.2332 0.1538 0.1053 0.1501 0.2827 0.4091 0.2493 0.1594 0.1045 0.1509 0.3181
Yang et al. [48] (2021) 0.3585 0.2266 0.1550 0.1021 0.1425 0.2788 0.3833 0.2409 0.1598 0.1078 0.1457 0.3293
XProNet [42] (2022) 0.3532 0.2212 0.1498 0.1052 0.1415 0.2811 0.4114 0.2502 0.1598 0.1045 0.1457 0.3240
ITHN [39] (2023) 0.3623 0.2128 0.1402 0.0992 0.1488 0.2622 0.2661 0.1516 0.0976 0.0663 0.1561 0.2617
M2KT [47] (2023) 0.3661 0.2192 0.1465 0.1044 0.1528 0.2673 0.2559 0.1381 0.0819 0.0523 0.1468 0.2439
@@ 4.1 Experiment Setting
Datasets. We conduct our experiments on two main publicly available chest X-ray datasets, i.e. MIMIC-CXR [20] and OpenI [11] datasets. MIMIC-CXR dataset is the largest public dataset for chest radiographs, including 473,057 chest X-ray images and 206,563 reports from 63,478 patients. We follow the official splits, i.e. 368,960 instances for training, 2,991 for validation, and 5,159 for testing. OpenI is a chest X-ray dataset with 3,684 report-image pairs, including 2,912 pairs for training and 772 pairs for testing. Different from existing methods [7, 8, 48, 47], we adopt both finding and impression sections as medical reports and remove the reports with less than 3 tokens for two datasets. There is no overlap of patients among different subsets. The CXR images with various sizes are all resized to $256\times 256$ pixels.
Implementation Details We adopt the PVT-medium [43] pre-trained with the disease classification task as the image encoder with AdaPatch in stage 2, 3 and 4. The AdaMatch is optimized with LAMB optimizer [51] and cosine learning rate scheduler [29]. The learning rate is 6e-3. The maximum training epoch is 15 and the per GPU batch size is 112. In AdaMatch-Cyclic, we adopt the VQ-GAN [13] models pre-trained on the MIMIC-CXR and OpenI datasets, respectively, and the dolly-v2-3b [9] model as pre-trained LLM. LLM has 5,1845 token types with 5,0821 token types for text tokens and 1,024 token types for image tokens. LLM is optimized by the AdamW [30] optimizer with a learning rate of 5e-6 and per GPU batch size of 24. The training epoch is 5. 8 Nvidia A100 40GB GPUs are used for all experiments. For more details, see supplementary materials.
Evaluation Metrics We evaluate the CXR-to-report generation performance with BLEU [33], METEOR [12], and ROUGE-L [27]. The FID [15] and NIQE [31] are utilized to evaluate report-to-CXR generation. The performance of CXR-to-report and report-to-CXR retrieval is assessed through the exact report in the top K retrieved reports for a given CXR image (R@K, K=$\left\{1,5,10\right\}$).
@@ 4.2.1 CXR-to-Report Generation
We verify the effectiveness of AdaMatch-Cyclic on the MIMIC-CXR and OpenI datasets for report generation. Since previous methods [7, 8, 48, 42, 39, 47] are mainly trained on the medical reports with the finding section, we reimplement previous methods on datasets with both finding and impression sections. In Table 1, our AdaMatch-Cyclic substantially surpasses the M2KT [47] by a significant margin with the BLEU-1 of 0.0132 in the MIMIC-CXR dataset. Moreover, we evaluate the OpenI dataset with the model trained on the MIMIC-CXR dataset. In terms of the zero-shot report generation, our AdaMatch-Cyclic achieves the best performance with the BLEU-1, METEOR, ROUGE-L of 0.4161, 0.1621, and 0.3656. To further compare with GPT-4V[49], we randomly select 50 cases from the test set of the MIMIC-CXR dataset. In Table 2, our AdaMatch-Cyclic significantly outperforms the GPT-4V by 0.1479 and 0.0892 of BLEU-1 and BLEU-4, indicating our superiority to existing methods.
Table 2: CXR-to-report generation performance in comparison to GPT-4V on 50 selected cases of the MIMIC-CXR dataset.
Methods BLEU-1 BLEU-2 BLEU-3 BLEU-4
R2Gen[7](2020) 0.3513 0.2174 0.1447 0.1025
R2GenCMN[8](2021) 0.3587 0.2200 0.1436 0.0969
Yang et al. [48](2021) 0.3596 0.2218 0.1481 0.1026
XProNet[42](2022) 0.3356 0.2071 0.1374 0.0941
ITHN[39](2023) 0.3301 0.1839 0.1121 0.0723
M2KT[47](2023) 0.3626 0.2123 0.1391 0.0957
GPT-4V[49](2023) 0.2275 0.0878 0.0378 0.0166
AdaMatch-Cyclic 0.3754 0.2303 0.1520 0.1058
Table 3: Report-to-CXR generation performance in comparison to existing methods on MIMIC-CXR and OpenI datasets.
MIMIC-CXR OpenI
[…]
Stable diffusion [37] 9.2334 3.7894 8.2946 6.3496
Chambon et al. [3] 8.2758 3.8871 5.8557 4.6534
RoentGen [2] 9.5411 3.8834 6.5675 4.9085
UniXGen [22] 6.7212 3.7125 11.9890 4.6610
LLM-CXR [23] 2.1788 3.5969 1.6597 3.8206
AdaMatch-Cyclic 1.0916 3.3931 1.5938 3.3096
To qualitatively evaluate report generation performance, we show two cases of generated reports in Fig. 5. In comparison to M2KT [47] and R2GenCMN [8], our AdaMatch-Cyclic can obtain the related keywords (e.g. scarring and atelectasis) and correctly describes the ‘osseous abnormality’ and ‘clips in the right upper quadrant of the abdomen’ in the first case, and point out ‘mild granulomatous’ in the second case, suggesting that our method can generate medical reports with higher quality than previous methods.
4.2.2 Report-to-CXR Generation
To quantitatively assess the performance, we compare our method with the text-to-image generation method [37] and some report-to-CXR generation methods [3, 2, 22, 23]. As listed in Table 3, AdaMatch-Cyclic achieves the best performance with 1.0916 and 1.5938 of FID scores on the MIMIC-CXR and OpenI datasets, respectively, indicating the superiority of our method over previous methods. To compare the high-level feature distribution of CXR images generated by different methods, we randomly select 1,000 cases from the test set, and apply the t-SNE visualization to the real and synthetic CXR images on the MIMIC-CXR dataset. As depicted in Fig. 6, synthetic CXR images generated by existing methods and real ones differ significantly from each other, while CXR images synthesized by AdaMatch-Cyclic almost overlap with the real CXR images. It implies that AdaMatch-Cyclic is more effective in translating medical reports to realistic CXR images compared to existing methods.
@@ 4.2.2 Report-to-CXR Generation
Figure 7: The generated chest X-ray images of the MIMIC-CXR (1st row) and OpenI (2nd row) datasets with highlighted regions.
In Fig. 7, we show the generated CXR images of the MIMIC-CXR and OpenI datasets. In the first example, our AdaMatch-Cyclic can synthesize the ‘left and right pleural effusions’ as highlighted in the CXR images, while other methods [3, 2, 37] cannot generate the CXR images with ‘pleural effusions’. It suggests that our method can generate more realistic CXR images than existing methods according to the input report.
4.2.3 CXR-Report Retrieval
To prove the superiority of AdaMatch, we evaluate the CXR-to-report and report-to-CXR retrieval performance on the MIMIC-CXR datasets in comparison with existing methods. As listed in Table 4, our AdaMatch achieves the best retrieval performance, surpassing the LIMITR [10] by a large margin with R@1 of 11.77% and 13.48% for CXR-to-report and report-to-CXR retrieval, respectively, implying that our method is more effective in extracting distinct semantic features for CXR images and medical reports and is able to align their features correctly.
Table 4: CXR-report retrieval performance (%) in comparison to existing methods on the MIMIC-CXR dataset.
@@ 4.2.3 CXR-Report Retrieval
Fang et al. [14] (2015) 18.60 43.10 56.10 18.13 43.20 55.97
Chauhan et al. [4] (2020) 5.37 19.43 30.73 5.40 20.23 30.23
ConVIRT [53] (2020) 30.10 53.90 63.80 29.20 54.70 64.40
GLoRIA [16] (2021) 30.30 57.50 66.50 24.00 51.80 62.80
JoImTeR-Net [18] (2021) 18.93 46.20 58.67 19.07 45.27 58.50
MGCA [41] (2022) 25.80 51.90 62.10 27.90 51.20 61.60
LIMITR [10] (2023) 39.70 63.20 71.70 37.70 62.10 71.30
Motor [26] (2023) 10.96 31.93 42.90 12.00 33.10 44.32
@@
4.3 Ablation Study
Effectiveness of AdaPatch. To prove the effectiveness of AdaPatch, we compare the CXR-report retrieval performance of AdaMatch model with or without AdaPatch. In the AdaMatch model without AdaPatch, we use PVT-medium [43] as the image encoder. In Table 7, the AdaPatch module remarkably improves the R@1 by 2.70% and 2.46% for CXR-to-report and report-to-CXR retrieval, respectively, indicating the effectiveness of our AdaPatch.
Image Encoder with Different Stages. In AdaMatch, the image encoder consists of four stages, and the stage 2,3 and 4 include an AdaPatch module, respectively. To investigate the effectiveness of each stage for CXR-report alignment, we use patch embeddings from different stages to compute the CXR-report similarity for CXR-report retrieval task. In Table 6, the image encoder with three stages achieves the best retrieval performance with the R@1 of 51.47%, implying that the features of the 3rd stage are more beneficial for patch-word matching.
@@ 4.3 Ablation Study
Visualization of Adaptive Patches. To visualize adaptive patches in AdaPatch, we display several examples from the MIMIC-CXR datasets. In Fig. 8, each example includes the text, its text tokens, and the corresponding CXR image with the marked adaptive patches. The adaptive patches are marked by the bounding boxes with different colors. Each bounding box includes an arrow, which shows the shift of center from the fixed patch to the adaptive patch. In the first example, adaptive patches cover the pacemaker and its wire. Meanwhile, adaptive patches of the second example show the correct position of pleural effusions in the CXR image. These indicate that AdaPatch can correctly localize the region relevant to input text tokens.
Figure 8: Visualization of texts and the corresponding adaptive patches. The boxes and arrows in different colors show adaptive patches and their shift of center from fixed patches.

### v5

@@ 3.2.1 CXR-to-Report Generation
In CXR-to-report generation, we use AdaMatch to match a predefined textual codebook with the CXR image, providing keywords to guide LLM.
Building textual codebook. Specifically, we first use a pre-trained BioEN Raza et al. (2022) to extract the related entities from medical reports in the training set, where the related entities are divided into four entity groups, i.e., biological structure, detailed description, disease disorder, and sign symptom. Next, we compute the frequency of each entity and pick top $\kappa_{0}$ entities for each entity group as keywords in the textual codebook.
Keywords Extraction. With the textual codebook, we employ the frozen AdaMatch model to match keywords from the textual codebook with the adaptive patches of CXR images, and obtain a patch-word similarity matrix $s^{P}\in\mathbb{R}^{M\times N}$ between the keyword tokens and adaptive patch embeddings, where $M$ and $N$ denote the number of keyword tokens and adaptive patches. To extract the most matched keywords for each adaptive patch, we rank the patch-word similarity along the dimension of keyword tokens, obtain the top $\kappa_{1}$ patch-word similarities for each adaptive patch, and extract corresponding keywords $k^{T}$. The matched keywords can explain potential lesions in each adaptive patch, to assist LLM in generating medical reports.
Figure 3: The example of instruction data for CXR-to-report generation.
Instruction tuning LLM. After keywords extraction, we use a frozen VQ-GAN Esser et al. (2021) encoder $E$ to encode the input CXR image $x^{I}$ to the quantized image latent vectors as image tokens $E(x^{I})$. For CXR-to-report generation, we adopt the dolly-v2-3b Conover et al. (2023) model as the pre-trained LLM, and convert CXR-to-report generation dataset in instruction-following format. An example of instruction data for CXR-to-report generation is depicted in Fig. 3, with the instruction, input (the image tokens of the input CXR image), keywords (extracted by AdaMatch), and response parts (the ground-truth medical report). During training, the LLM learns to generate the hidden response part in an autoregressive manner.
By adopting AdaMatch-Cyclic, we can locate potential lesions and interpret them with keywords to enhance explainability.
@@ 3.2.2 Report-to-CXR Generation
In report-to-CXR generation, AdaMatch matches a predefined visual codebook with generated reports $x^{T^{\prime}}$, providing keypatches as guidance for the LLM to synthesize CXR images, where keypatches are important image patches related to the reports.
Building visual codebook. Concretely, we first construct a visual codebook with the most common adaptive patches as keypatches. To collect the most common adaptive patches for CXR images in the training set, AdaMatch matches the adaptive patches of CXR images with textual tokens of medical reports to obtain top $\kappa_{2}$ CXR-report pairs with the highest report-to-CXR similarities $s^{T}$. For each CXR-report pair, we compute the word-patch maximum similarity $\max_{(1\leq n\leq N)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$ for each textual token, rank the word-patch maximum similarities, and extract the adaptive patches for top $\kappa_{3}$ similarities as keypatches in the visual codebook. Each keypatch includes its adaptive patch and the corresponding features.
Keypatches Extraction. With the visual codebook, the frozen AdaMatch matches the features of keypatches in the visual codebook with textual tokens of the generated report to acquire the word-patch similarity matrix $s^{W}\in\mathbb{R}^{(\kappa_{2}\times\kappa_{3})\times K}$, where $K$ is the number of textual tokens. To obtain keypatches related to the generated report, we rank the word-patch similarity along the dimension of keypatches, obtain the top $\kappa_{4}$ word-patch similarity for each textual token, and extract the features of corresponding keypatches $k^{I}$.
Table 1: Comparison of CXR-to-report generation performance on the MIMIC-CXR and the OpenI datasets.
MIMIC-CXR OpenI
Methods B-1 B-2 B-3 B-4 M R-L B-1 B-2 B-3 B-4 M R-L
R2Gen 0.3553 0.2232 0.1523 0.1038 0.1412 0.2784 0.3992 0.2407 0.1518 0.0973 0.1390 0.3052
R2GenCMN 0.3719 0.2332 0.1538 0.1053 0.1501 0.2827 0.4091 0.2493 0.1594 0.1045 0.1509 0.3181
Joint-TriNet 0.3585 0.2266 0.1550 0.1021 0.1425 0.2788 0.3833 0.2409 0.1598 0.1078 0.1457 0.3293
XProNet 0.3532 0.2212 0.1498 0.1052 0.1415 0.2811 0.4114 0.2502 0.1598 0.1045 0.1457 0.3240
ITHN 0.3623 0.2128 0.1402 0.0992 0.1488 0.2622 0.2661 0.1516 0.0976 0.0663 0.1561 0.2617
M2KT 0.3661 0.2192 0.1465 0.1044 0.1528 0.2673 0.2559 0.1381 0.0819 0.0523 0.1468 0.2439
Instruction tuning LLM. After keypatches extraction, we use a frozen VQ-GAN encoder to convert the matched keypatches $k^{I}$ into image tokens $E(k^{I})$, and feed the LLM with the instruction, generated report, and image tokens of keypatches in the instruction-following format, as shown in Fig. 4. Then, LLM predicts image tokens, decoded by the VQ-GAN into the generated CXR image $x^{I^{\prime}}$.
Figure 4: The example of instruction data for report-to-CXR generation.
@@ 3.2.3 Overall Objective
To optimize AdaMatch-Cyclic, we use the standard language modeling objective for both CXR-to-report and report-to-CXR generation. For CXR-to-report generation, LLM receives the instruction $I^{t}$, image tokens of input CXR image $E(x^{I})$, and matched keywords $k^{T}$, aiming to generate the ground-truth medical report of response part autoregressively. We compute the conditional probability of the $k$-th token, $P(u_{k})=\mathcal{LLM}(I^{t},E(x^{I}),k^{T})$, where $k$ is the token index after the response key (### Response:). The report generation loss for the response area is calculated by, $L_{text}=\sum_{i=k}^{n}-\log P(u_{i}|u_{1},u_{2},\cdots,u_{i-1})$, where $[u_{1},u_{2},\cdots,u_{i-1}]$ denotes the tokenized texts before the response part, and $n$ represents the maximum length of output tokens. Similarly, for report-to-CXR generation, we feed the LLM with the instruction $I^{v}$, generated report $x^{T^{\prime}}$, and image tokens of keypatches $E(k^{I})$, and obtain the conditional probability of $k$-th token $w_{k}$, $P(w_{k})=\mathcal{LLM}(I^{v},x^{T^{\prime}},E(k^{I}))$. The CXR generation loss is defined as: $L_{img}=\sum_{i=k}^{n}-\log P(w_{i}|w_{1},w_{2},\cdots,w_{i-1})$. With $L_{text}$ and $L_{img}$, LLM can implement bidirectional CXR and report generation according to instructions.
Figure 5: Qualitative comparison of CXR-to-report generation on the MIMIC-CXR (1st row) and the OpenI (2nd row) datasets, highlighting similar meanings in colored text. The keywords are obtained from AdaMatch.
@@ 4.1 Experiment Setting
Datasets. We experiment on two main publicly available chest X-ray datasets, i.e. MIMIC-CXR Johnson et al. (2019) and OpenI Demner-Fushman et al. (2016) datasets. MIMIC-CXR dataset comprises 473,057 images and 206,563 reports from 63,478 patients. We use the official splits, i.e. 368,960 for training, 2,991 for validation, and 5,159 for testing. OpenI contains 3,684 report-image pairs with 2,912 for training and 772 for testing. Unlike prior methods, we utilize both finding and impression sections as medical reports.
Table 2: Performance of CXR-to-report generation compared to GPT-4V on 50 selected MIMIC-CXR cases.
Methods B-1 B-2 B-3 B-4
R2Gen 0.3513 0.2174 0.1447 0.1025
R2GenCMN 0.3587 0.2200 0.1436 0.0969
Joint-TriNet 0.3596 0.2218 0.1481 0.1026
XProNet 0.3356 0.2071 0.1374 0.0941
ITHN 0.3301 0.1839 0.1121 0.0723
M2KT 0.3626 0.2123 0.1391 0.0957
GPT-4V 0.2275 0.0878 0.0378 0.0166
AdaMatch-Cyclic 0.3754 0.2303 0.1520 0.1058
Table 3: Comparison of report-to-CXR generation performance on the MIMIC-CXR and the OpenI datasets.
MIMIC-CXR OpenI
[…]
Stable diffusion 9.2334 3.7894 8.2946 6.3496
Adapting-Med 8.2758 3.8871 5.8557 4.6534
RoentGen 9.5411 3.8834 6.5675 4.9085
UniXGen 6.7212 3.7125 11.9890 4.6610
LLM-CXR 2.1788 3.5969 1.6597 3.8206
AdaMatch-Cyclic 1.0916 3.3931 1.5938 3.3096
Implementation Details. In the AdaMatch-Cyclic model, we first train the AdaMatch model and then use the frozen AdaMatch to train LLM. We adopt VQ-GAN Esser et al. (2021) models pre-trained on the MIMIC-CXR and OpenI datasets, respectively, and the dolly-v2-3b Conover et al. (2023) model as pre-trained LLM. The source code will be released. Please see appendix C for more details.
Evaluation Metrics. We assess CXR-to-report generation using BLEU (B), METEOR (M), and ROUGE-L (R-L) Chen et al. (2020b), and report-to-CXR generation using FID Heusel et al. (2017) and NIQE Mittal et al. (2012). The retrieval performance between CXR and report is assessed through the exact report in the top K retrieved reports for a given CXR image (R@K, K=$\left\{1,5,10\right\}$).
@@ 4.2.1 CXR-to-Report Generation
We compare AdaMatch-Cyclic with current CXR-to-report generation methods on the MIMIC-CXR and the OpenI datasets, including R2Gen Chen et al. (2020b), R2GenCMN Chen et al. (2021a), Joint-TriNet Yang et al. (2021), XProNet Wang et al. (2022b), ITHN Voutharoja et al. (2023), and M2KT Yang et al. (2023). Since these methods are mainly trained on the finding section, we reimplement them on both the finding and impression sections. In Table 1, AdaMatch-Cyclic achieves the best performance of 0.3793 in BLEU-1 on the MIMIC-CXR dataset, with superior generalization to OpenI dataset using the same model. Compared to GPT-4V, AdaMatch-Cyclic significantly outperforms in BLEU-1 by 0.1479, as listed in Table 2.
@@
Figure 7: Generated CXR images of the MIMIC-CXR (1st row) and OpenI (2nd row) datasets with highlighted regions.
Table 4: Comparison of CXR-report retrieval performance (%) on MIMIC-CXR dataset.
@@
Fang et al. (2015) 18.60 43.10 56.10 18.13 43.20 55.97
Chauhan et al. (2020) 5.37 19.43 30.73 5.40 20.23 30.23
ConVIRT Zhang et al. (2022) 30.10 53.90 63.80 29.20 54.70 64.40
GLoRIA Huang et al. (2021) 30.30 57.50 66.50 24.00 51.80 62.80
JoImTeR-Net Ji et al. (2021) 18.93 46.20 58.67 19.07 45.27 58.50
MGCA Wang et al. (2022a) 25.80 51.90 62.10 27.90 51.20 61.60
LIMITR Dawidowicz et al. (2023) 39.70 63.20 71.70 37.70 62.10 71.30
Motor Lin et al. (2023) 10.96 31.93 42.90 12.00 33.10 44.32
@@
Table 6: Effectiveness of report-to-CXR and CXR-to-report generation tasks.
[…]
0.3542 0.1973 0.1190 0.0758 0.1329 0.2392 - -
- - - - - - 1.7128 4.0391
0.3793 0.2346 0.1540 0.1060 0.1625 0.2813 1.0916 3.3931
Table 7: Comparison of AdaPatch and other adaptive vision models.
CXR-to-Report Report-to-CXR
Models R@1 R@5 R@10 R@1 R@5 R@10
A-ViT 49.78 83.14 92.77 49.49 83.19 92.94
AdaViT 50.14 84.52 93.20 50.54 84.10 93.83
AdaPatch 51.47 86.19 94.77 51.18 86.46 94.60
@@
Figure 8: Visualization of texts and its adaptive patches. Boxes and arrows in different colors show adaptive patches and their center shifts from fixed patches.
In Fig. 5, we compare AdaMatch-Cyclic’s performance with M2KT and R2GenCMN. AdaMatch-Cyclic accurately captures relevant keywords like ‘scarring’, and describes ‘osseous abnormality’ and ‘clips in the right upper quadrant of the abdomen’ in the first case, showcasing superior report quality over previous methods.
4.2.2 Report-to-CXR Generation
We quantitatively compare our method with text-to-image generation method like Stable Diffusion Rombach et al. (2022) and report-to-CXR generation methods, such as Adapting-Med Chambon et al. (2022b), RoentGen Chambon et al. (2022a), UniXGen Lee et al. (2023a), and LLM-CXR Lee et al. (2023b). In Table 3, AdaMatch-Cyclic achieves the highest FID scores on both datasets, showing its superior effectiveness in generating CXR. To compare the high-level feature distribution of CXR images generated by different methods, we randomly select 1,000 cases from test set, and apply t-SNE visualization to the real and synthetic CXR images on the MIMIC-CXR dataset. In Fig. 6, while existing methods’ synthetic CXR images differ from real ones, AdaMatch-Cyclic’s almost overlap with real ones, indicating its superiority in report-to-CXR generation.
In Fig. 7, we display generated CXR images from the MIMIC-CXR and OpenI datasets. In the first example, AdaMatch-Cyclic accurately synthesizes ‘left and right pleural effusions’, while other methods fail to generate such features. This suggests the superior ability of our method to produce realistic CXR images based on input reports.
4.2.3 CXR-Report Retrieval
To demonstrate the superiority of AdaMatch, we evaluate the CXR-to-report and report-to-CXR retrieval performance on the MIMIC-CXR dataset compared to existing methods. We utilize the AdaMatch to compute the similarities between CXR images and reports, rank the similarities, and obtain the retrieval results. As shown in Table 4, AdaMatch significantly outperforms LIMITR with R@1 of 11.77% and 13.48% for CXR-to-report and report-to-CXR retrieval. This indicates the superior ability of our method to extract distinct semantic features and align them accurately between CXR images and medical reports.
In Table 5, to ablate AdaPatch, we remove the layers to predict offset and scale. We then aligned grid features from the last stage of the pyramid vision Transformer with textual tokens to compute the image-text similarity for retrieval tasks.
4.3 Ablation Study
Effectiveness of AdaPatch. We evaluate the effectiveness of AdaPatch by comparing the CXR-report retrieval performance of AdaMatch with and without AdaPatch. To ablate AdaPatch, we remove the layers to predict offset and scale, and compute the similarity between grid features of the last stage and textual tokens for retrieval tasks. In Table 5, AdaPatch significantly improves R@1 by approximately 3% for both retrieval directions, highlighting its effectiveness. To demonstrate the optimality of the AdaPatch design, we compare AdaPatch with other adaptive vision models, such as A-ViT Yin et al. (2022) and AdaViT Meng et al. (2022), in the CXR-report retrieval task. As shown in Table 7, AdaPatch exhibits superior retrieval performance compared with other methods.
Effectiveness of Cyclic Generation. To validate the necessity of report-to-CXR and CXR-to-report generation tasks, we remove each task and evaluate their performance, as listed in Table 6. When ablating report-to-CXR generation task, the performance of CXR-to-report generation decreases substantially compared to our method, and vice versa. These suggest that the CXR-to-report and report-to-CXR generation can indeed benefit each other, and their interaction is crucial for the overall performance of the model.
Image Encoder with Different Stages. In AdaMatch, the image encoder comprises several stages, with AdaPatch modules in the stages. To assess the effectiveness of the stage number, we compare the retrieval performance of AdaMatch with two, three, or four stages. Table 8 shows that the three-stage image encoder achieves the highest R@1 of 51.47%, indicating that the features from the 3rd stage are most beneficial for CXR-report retrieval.
Visualization of Adaptive Patches. We visualize adaptive patches in AdaPatch of some examples from the MIMIC-CXR dataset in Fig. 8. Each example includes text, its tokens, and the corresponding CXR image with marked adaptive patches. The patches are represented by colored bounding boxes with arrows indicating the center shift from the fixed patch. In the first example, adaptive patches cover the pacemaker and wire, and the second one can find pleural effusions, implying the capacity of AdaPatch to localize regions relevant to input text tokens accurately.